55 students are selected at random from an engineering class and classified as

male or female. List the elements of sample space S1 using the letter M for male

F for female. Define a second sample space S2 Where the elements represent the

number of females selected.


1
Expert's answer
2021-09-17T03:44:53-0400

"S_1" = {(55 "\times" M and 0 "\times" F),

(55-1 "\times" M and 1 "\times" F),

(55-2 "\times" M and 2 "\times" F),…

(0 "\times" M and 55 "\times" F)}


"S_2" = {0,1,2,…,53,54,55}